# Quaternionic Pole Placement via Companion Forms and the Ackermann Formula

### v1

@@ VIII Conclusion
This note developed a quaternionic framework for state-feedback pole placement. We (i) formalized controllable companion forms over $\mathbb{H}$ and their companion polynomials, which right-annihilate the companion matrix, and (ii) gave two practical designs: coefficient matching in controllable coordinates and a coordinate-free Ackermann construction valid for real (central) targets. The method keeps the convenience of field-based workflows while addressing quaternionic specifics—companion polynomial instead of characteristic/minimal, right-eigenvalue similarity classes, and order-sensitive polynomial evaluation.
Although we focused on single-input systems, extension to multi-input appears straightforward via block companion/Brunovsky forms and the same central-polynomial calculus—mainly algebraic bookkeeping. Future work will focus on observer duals, robustness, the numerical conditioning of quaternionic companion computations, and output-feedback designs.

### v2

@@ VIII Conclusion
This note developed a quaternionic framework for single-input state-feedback pole placement. We formalized controllable companion forms over $\mathbb{H}$ and their companion polynomials satisfying $a(A_{c})_{\mathscr{r}}=0$, and presented two designs: coefficient matching in companion coordinates and a coordinate-free Ackermann gain for real (central) targets. The approach mirrors classical workflows while making quaternionic specifics explicit (similarity-class poles and order-sensitive evaluation).
Multi-input extensions should follow via block companion/Brunovsky forms with the same central-polynomial calculus. Future work includes observer duals, robustness, numerical conditioning, and output-feedback design.